OpenStudy (anonymous):

A vending machine offers 8 different drinks. One day, 6 employees each purchased a drink from the vending machine. Find the probability that at least 2 employees purchased the same drink. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

hmm

OpenStudy (perl):

each employee has 8 choices. So total cases = 8^6. each employee has different drink: Number of cases = 8*7*6*5*4*3. for at least two have same drink = total cases - when each have different drink = 8^6 - 8*7*6*5*4*3. probability = (8^6 - 8*7*6*5*4*3) / (8^6).

OpenStudy (perl):

probability two or more have same drink = 1 - probability no two have same drink

OpenStudy (anonymous):

yeah i understand that but when i plugg it in calculatior i get wrong answer

OpenStudy (anonymous):

@perl are those times signals

OpenStudy (perl):

yes

OpenStudy (perl):

the answer is 3781/4096 = .9230957

OpenStudy (anonymous):

i was plugging in wrong hmm no wonder ahh thanks perl your so smart :)

OpenStudy (anonymous):

@perl is it rounded to the nearest hundreth?

OpenStudy (anonymous):

.92 :) nvm